# Gentle Object Retraction in Dense Clutter Using Multimodal Force Sensing and Imitation Learning

### v1

@@ VII Conclusion & Future Work
In this work, we demonstrate the ability to use imitation learning to train autonomous policies for gently navigating and retracting objects from extremely variable, densely cluttered lateral access environments. We find that in contrast to a baseline policy without access to wrench and tactile information, any policy with access to one or both force modalities demonstrates an improvement in overall success rate, a reduction in excess force impulse failures, and an improvement to the average success time. The fully force-informed policy shows the best performance, with an improvement of 8080% over the baseline.
@@
Additional improvements to the system could involve adding an external visual stream that captures an isometric view of the scene or using a pre-trained segmentation network to identify arbitrary objects rather than only red targets. The network architecture could be improved by fine tuning policies with reinforcement learning as in [36], incorporating more reactive diffusion policy architectures as in [37], or building a custom tactile encoder from play data as in [5]. The proposed imitation learning framework could also be compared against a system that implements a model-based, low-level force-informed controller as in [38] integrated with learned motion planning using vision and proprioception.
More broadly, this work represents a call to move beyond robot interaction in mere visual clutter and to embrace the degree of physical clutter encountered on shelves and in cabinets in homes and commercial settings. By learning from the hardware, sensing, and reactive strategies that humans use on a daily basis, we can enable robots to perform robust, contact-rich manipulation in unstructured environments.

### v2

@@ VII Conclusion & Future Work
In this work, we demonstrate the ability to use imitation learning to train autonomous policies for gently navigating and retracting objects from extremely variable, densely cluttered lateral access environments. We find that in contrast to a baseline policy without access to wrench and tactile information, any policy with access to one or both force modalities demonstrates an improvement in overall success rate, a reduction in excess force impulse failures, and an improvement to the average success time. The wrench + tactile policy shows the best performance, with an improvement of 8080% over the baseline.
@@
Furthermore, an additional study of the system’s robustness to out of distribution objects for both obstacles and targets could prove useful to further investigate the importance of tactile and wrench information in these scenarios. Using a pre-trained segmentation network may be necessary to identify arbitrary objects rather than only red targets. To alleviate the limitation that the force thresholds used in this work are task specific, adaptive gentleness could be implemented, perhaps with control barrier function approaches [42, 43]. The network architecture could be improved by fine tuning policies with reinforcement learning as in [44], incorporating more reactive diffusion policy architectures as in [45], or building a custom tactile encoder from play data as in [5]. The proposed imitation learning framework could also be compared against a system that implements a model-based, low-level force-informed controller as in [46] integrated with learned motion planning using vision and proprioception. Investigating whether the benefits of contact force sensing modalities reported in this work are echoed in alternative control architectures would be an especially valuable future direction.
More broadly, this work represents a call to move beyond robot interaction in mere visual clutter and to embrace the degree of physical clutter encountered on shelves and in cabinets in homes and commercial settings. By learning from the embodied, sensing, and reactive strategies that humans use on a daily basis, we can enable robots to perform robust, contact-rich manipulation in unstructured environments.